# Perturbation-based Effect Measures for Compositional Data

### v4

@@ Acknowledgements
S4.1.1 Nonparametric estimation for directional effects
We begin with the nonparametric estimation of a directional perturbation effect $\tau$ corresponding to Theorem S1 (a). A high-level overview of this estimation was given in Section 3.2.2. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, a differentiable regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$, a score estimation method $\widehat{\rho}$ for the score $(\ell,w)\mapsto\partial_{\ell}\log(p(\ell\,|\,w))$, where $p$ denotes the conditional density of $L$ given $W$ with corresponding nuisance function estimation method $\widehat{\eta}$ (in our experiments, we assume a location-scale model and thus the nuisance functions are estimators of the conditional mean and variance of $L$ given $W$, see Section S4.6). We can then compute the estimator $\widehat{\tau}$ of $\tau$ and its corresponding variance estimate $\widehat{\sigma}_{\tau}^{2}$ as described in Algorithm S1 using a user-specified number of splits $K$ in the cross-validation.
@@
S4.1.2 Nonparametric estimation for binary effects
We now present the algorithm and conditions required for the nonparametric estimation of the binary perturbation effect $\lambda$ corresponding to Theorem S1 (b). A high-level overview of this estimation was provided in Section 3.2.1. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, a regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and propensity score estimator $\widehat{\pi}$ for the propensity score $w\mapsto\mathbb{P}(L=1\,|\,W=w)$. We can then compute the estimator $\widehat{\lambda}$ of $\lambda$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\lambda}$ as described in Algorithm S2 using a user-specified number of splits $K$ in the cross-validation.
@@
S4.1.3 Partially linear model for binary/directional effects
Finally, we present the algorithm and conditions required for the estimation of the perturbation effects under a partially linear model assumption corresponding to Theorem S1 (c). A high-level overview of this estimation was provided in Section 3.2.3. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, regression methods $\widehat{m}$ and $\widehat{g}$ for the conditional expectations $w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $w\mapsto\mathbb{E}[L\,|\,W=w]$, respectively. We can then compute the estimator $\widehat{\theta}$ of $\theta$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\theta}$ as described in Algorithm S3 using a user-specified number of splits $K$ in the cross-validation.
@@
S4.3 Robustness of nuisance function estimation for different estimators
@@
S4.3.1 Estimation of nuisance functions with continuous $L$
To illustrate the difficulty of estimating the nuisance functions $\partial_{\ell}f$ and $\rho$ (see Section 3.2.2) compared to $m$ and $g$ (see Section 3.2.3) we estimate the functions on simulated data and compare the estimates to the truth. We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ and $\xi$ are standard Gaussian independent of each other and from $W$ and
@@
S4.3.2 Estimation of nuisance functions with binary $L$
To illustrate the difficulty of estimating the inverse propensity score $\pi^{-1}$ (see Section 3.2.1) compared to $m$ (in this case equal to $\pi$) and $g$ (see Section 3.2.3) we estimate the functions on simulated data and compare the estimates to the truth. We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ is standard Gaussian independent of $W$, $L$ conditioned on $W$ is generated as a Bernoulli random variable with success probability $\pi(W):=0.95\mathrm{expit}(\mathrm{logit}(W^{1}))+0.05$ and
@@
S4.3.3 Partially linear model estimate when partial linearity is violated
@@
S4.4 Computing directional perturbation effects in the presence of zero speeds
@@
S4.5 Estimating differentiable regression functions
@@
S4.6 Location-scale modelling of the score function
@@
Section S5.1 provides the details behind the experiments in Table S1 in Section 1. Section S5.2 contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional. Section S5.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 4.1.1. Section S5.4 includes figures visualizing the semi-synthetic dataset used in Section 4.1.1. Section S5.6 provides a detailed description of the non-compositional background variables that are controlled for in Section 4.1.2. Section S5.7 contains additional plots illustrating results from the microbiome data analysis in Section 4.2.
@@
  • NPM: The one-step corrected nonparametric estimators including cross-fitting, i.e., (10) and (11). See Algorithms S1 and S2, with $K=2$, for the detailed binary and directional effect estimators, respectively.
@@
  • plugin: The naive plug-in versions of the NPM estimates. In the binary effect case, the estimator in (10) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (11) without the term involving $\widehat{\rho}$.
@@
S5.7 Additional plots from gut microbiome analysis in Section 4.2

### v5

@@ Acknowledgements
S4.2.1 Nonparametric estimation for directional effects
We begin with the nonparametric estimation of a directional perturbation effect $\tau$ corresponding to Theorem S1 (a). A high-level overview of this estimation was given in Section S4.1.2. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, a differentiable regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$, a score estimation method $\widehat{\rho}$ for the score $(\ell,w)\mapsto\partial_{\ell}\log(p(\ell\,|\,w))$, where $p$ denotes the conditional density of $L$ given $W$ with corresponding nuisance function estimation method $\widehat{\eta}$ (in our experiments, we assume a location-scale model and thus the nuisance functions are estimators of the conditional mean and variance of $L$ given $W$, see Section S4.7). We can then compute the estimator $\widehat{\tau}$ of $\tau$ and its corresponding variance estimate $\widehat{\sigma}_{\tau}^{2}$ as described in Algorithm S1 using a user-specified number of splits $K$ in the cross-validation.
@@
S4.2.2 Nonparametric estimation for binary effects
We now present the algorithm and conditions required for the nonparametric estimation of the binary perturbation effect $\lambda$ corresponding to Theorem S1 (b). A high-level overview of this estimation was provided in Section S4.1.1. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, a regression method $\widehat{f}$ for the conditional expectation $(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and propensity score estimator $\widehat{\pi}$ for the propensity score $w\mapsto\mathbb{P}(L=1\,|\,W=w)$. We can then compute the estimator $\widehat{\lambda}$ of $\lambda$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\lambda}$ as described in Algorithm S2 using a user-specified number of splits $K$ in the cross-validation.
@@
S4.2.3 Partially linear model for binary/directional effects
Finally, we present the algorithm and conditions required for the estimation of the perturbation effects under a partially linear model assumption corresponding to Theorem S1 (c). A high-level overview of this estimation was provided in Section 4.2. More formally, assume we have a dataset of $n$ i.i.d. copies of $(Y,L,W)$, regression methods $\widehat{m}$ and $\widehat{g}$ for the conditional expectations $w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $w\mapsto\mathbb{E}[L\,|\,W=w]$, respectively. We can then compute the estimator $\widehat{\theta}$ of $\theta$ and its corresponding variance estimate $\widehat{\sigma}^{2}_{\theta}$ as described in Algorithm S3 using a user-specified number of splits $K$ in the cross-validation.
@@
S4.4 Robustness of nuisance function estimation for different estimators
@@
S4.4.1 Estimation of nuisance functions with continuous $L$
To illustrate the difficulty of estimating the nuisance functions $\partial_{\ell}f$ and $\rho$ (see Section S4.1.2) compared to $m$ and $g$ (see Section 4.2) we estimate the functions on simulated data and compare the estimates to the truth. We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ and $\xi$ are standard Gaussian independent of each other and from $W$ and
@@
S4.4.2 Estimation of nuisance functions with binary $L$
To illustrate the difficulty of estimating the inverse propensity score $\pi^{-1}$ (see Section S4.1.1) compared to $m$ (in this case equal to $\pi$) and $g$ (see Section 4.2) we estimate the functions on simulated data and compare the estimates to the truth. We simulate $n=1000$ observations from the model where $W$ is uniformly distributed on $\Delta^{2}$, $\varepsilon$ is standard Gaussian independent of $W$, $L$ conditioned on $W$ is generated as a Bernoulli random variable with success probability $\pi(W):=0.95\mathrm{expit}(\mathrm{logit}(W^{1}))+0.05$ and
@@
S4.4.3 Partially linear model estimate when partial linearity is violated
@@
S4.5 Computing directional perturbation effects in the presence of zero speeds
@@
S4.6 Estimating differentiable regression functions
@@
S4.7 Location-scale modelling of the score function
@@
Section S5.1 provides the details behind the experiments in Table S1 in Section 1. Section S5.2 contains an investigation of the performance of the semiparametric estimators discussed in Section 4 when the conditioning variable $W$ is compositional. Section S5.3 provides a detailed description of how we generate the groups used for aggregation in the experiment described in Section 5.1.1. Section S5.4 includes figures visualizing the semi-synthetic dataset used in Section 5.1.1. Section S5.6 provides a detailed description of the non-compositional background variables that are controlled for in Section 5.1.2. Section S5.7 describes an additional experiment on the microbiome data analyzed in Section 5.2. Section S5.8 contains additional plots illustrating results from the microbiome data analysis in Section 5.2.
@@
  • NPM: The one-step corrected nonparametric estimators including cross-fitting, i.e., (S8) and (S9). See Algorithms S1 and S2, with $K=2$, for the detailed binary and directional effect estimators, respectively.
@@
  • plugin: The naive plug-in versions of the NPM estimates. In the binary effect case, the estimator in (S8) without the term involving $\widehat{\pi}$ and in the directional effect case, the estimator in (S9) without the term involving $\widehat{\rho}$.
@@
S5.7 Additional experiment on microbiome data in Section 5.2
In this section we describe an additional experiment on the microbiome data from Section 5.2. Our goal is to determinee how sensitive the proposed perturbation effect estimates are with respect to the employed nuisance function estimators. To this end, we consider three different regression methods: random forest (rf), neural network (mlp) and kernelized support vector regression using an RBF kernel (svr). Each of the regressions is combined with a grid-search cross-validation to automatically tune the hyperparameters of each method. For each of the three regression methods and all the species we then compute $\mathrm{CFI}_{\mathrm{mult}}$, $\mathrm{CFI}_{\mathrm{unit}}$ and $\mathrm{CKE}$ using the estimator based on a partially linear model and the $\mathrm{CKE}$ using the nonparametric estimator. As baselines, we additionally compute a nonparametric $R^{2}$ [Williamson et al., 2021] and permutation importance [Breiman, 2001]. Figure S11 shows the Spearman correlation between the estimates from the different regression methods.
Figure S11: Spearman correlations of variable importance measures computed on each feature using different regressions. On the left, these are computed on all $561$ features and on the right they are computed on the subset of features where at least one of the two methods deems the effect significant at a $5\%$ level. The $R^{2}$ and permutation methods do not come with significance tests and are therefore omitted from the right plot.
While the average perturbation effect estimates which are specifically tailored to the compositional setting are consistent across different regression methods, the permutation and $R^{2}$ measures are – unsurprisingly – not. The population version of the nonparametric $R^{2}$ is always 00, so the estimates are essentially mean zero random noise that depends on the estimator. For the permutation measure the inconsistency comes from the fact that the permuted data inherently leave the support of the simplex, making the resulting influence measure highly dependent on how the regression methods extrapolate. Unfortunately, due to a lack of alternatives, it is still common to use these types feature importance or influence measures in microbiome science [e.g., Marcos-Zambrano et al., 2021]. Furthermore, the results show that the nonparametric estimator of the CKE is less well-behaved than the other estimators based on the partially linear model. One reason for this could be that dividing by the propensity score estimator results in a more unstable estimate, particularly when the propensity score is poorly estimated.
S5.8 Additional plots from gut microbiome analysis in Section 5.2